# The Energy Cost of Reasoning: Analyzing Energy Usage in LLMs with Test-time Compute

### v1

@@ 4.2 In-depth Analysis
We conclude that parallel sampling with MV shows lower accuracy/energy than RT and that TTC has limited benefit on common-sense benchmarks. To this end, we provide a thorough analysis on RT using two reasoning benchmarks one from math (MATH500) and coding (LiveCodeBench) task each.
Figure 4: Energy vs Accuracy per length. Each line is represent a model and dot represents maximum sequence length.
@@
We investigate how the maximum output sequence length affects RT performance and associated computational costs. By incrementally varying the maximum sequence length by steps of 3277 tokens—one-tenth of the 32768 tokens—we assess the impact on accuracy-energy trade-off.
@@
Figure 5 illustrates the number of output tokens generated for correct and incorrect answers in both Base and RT models highlighting how token length relates to model reasoning.
Base models exhibit tightly clustered sequence lengths around 200 tokens for both correct and incorrect answers, with numerous outliers among incorrect responses. These models’ limited reasoning capabilities result in minimal token generation for answering, and outlier sequences typically represent unnecessary token generation rather than deeper reasoning. Increasing the model size reduces the number of outliers but maintains clustering near short sequence lengths, underscoring their restricted reasoning abilities.
@@
Secondly, for easier tasks, smaller models with RT can match the performance of larger models, resulting in significant energy savings. Specifically, deploying 14B and 1.5B models instead of a 32B model saves 11.55$\times$ and 1.48$\times$ energy for easy tasks in MATH500 and LiveCodeBench, respectively.
@@
Figure 7 depicts the QPS across various batch sizes for the two benchmarks using the highest-performing 32B models. Results indicate that larger batch sizes consistently improve throughput for both configurations. For Base, output sequences are relatively short, resulting in smaller KV cache footprints that under-utilize GPU memory BW. Although RT applied to MATH500 generates more tokens than Base, it still generates short enough sequences to under-utilize GPU memory BW. As a result, even at high batch sizes, the KV cache remains manageable, and throughput scales proportionally with batch size increases.
However, RT applied to the LiveCodeBench exhibit a notable saturation in throughput at a batch size of approximately 128. On average, RT generates around 6588 output tokens for LiveCodeBench, shifting inference into a predominantly decode-bound stage. During the decode stage, loading the large KV cache that scales linearly with batch size from GPU memory becomes the critical performance bottleneck, eventually saturating available GPU memory BW.
@@ 5 Conclusion and Discussion
Our results (Figure 4) demonstrate that smaller models with TTC can compete effectively against larger models, with distinct Pareto frontiers emerging for each task difficulty level. This insight supports the implementation of difficulty-aware evaluators [2024arXiv240407720S, park2024large, dutulescu2024hard, xu2024adaption] that dynamically route queries to the most energy-efficient model achieving optimal accuracy. Optimal energy consumption $E_{o}$ of an oracle evaluator and realistic energy consumption $E_{r}$ of an imperfect evaluator in practical settings employing a high-performing model as a fallback can be calculated as
where $p_{i}$ denotes the probability that the evaluator correctly predicts difficulty level $i$, $q$ represents the probability of incorrect predictions, and $e_{m_{D}}$ is the energy consumed by the largest, most accurate fallback model. Compared against the baseline approach ($E_{max}=Ne_{m_{D}}$), our analysis shows significant potential energy savings. For instance, using an oracle evaluator, LiveCodeBench and MATH500 benchmarks consumed only 1.75 kWh compared to baseline consumptions of 2.03 kWh and 2.20 kWh, respectively.

### v2

@@ 4.2 In-depth Analysis
We conclude that parallel sampling with MV shows lower accuracy/energy than RT and that TTC has limited benefit on common-sense benchmarks. To this end, we provide a thorough analysis on RT using one reasoning benchmark each from math (MATH500) and coding (LiveCodeBench) task.
Figure 4: Energy vs Accuracy per length. Dotted and solid lines represent Base and RT, respectively and each color represent different model size. Each of the ten dots on a line represent output sequence length limit starting from one-tenth of the maximum sequence length of a model to the maximum sequence length. Grey lines on the top represent the best models at the time of writing.
@@
We investigate how the maximum output sequence length affects RT performance and associated computational costs. We incrementally limit the output sequence length by steps of 3277 tokens which is one-tenth of the original maximum sequence length.
@@
Figure 5 illustrates the number of output tokens generated for correct and incorrect answers in both Base and RT models highlighting how sequence length relates to model reasoning.
Base models exhibit clustered sequence lengths for both correct and incorrect answers, with numerous outliers among incorrect responses. These models’ limited reasoning capabilities result in limited thinking process and outlier sequences typically represent unnecessary token generation rather than deeper reasoning. Increasing the model size reduces the number of outliers but maintains clustering near short sequence lengths, underscoring their restricted reasoning abilities.
@@
Figure 6: Accuracy vs energy per difficulty level. Each dot on a line represents different output sequence length limit identical to the previous Length Sweep analysis.
@@
[…]
@@
Figure 7 depicts the QPS across various batch sizes for the two benchmarks using the highest-performing 32B models. Results indicate that larger batch sizes consistently improve throughput for both configurations. For Base, output sequences are relatively short, resulting in smaller KV-cache footprints that under-utilize GPU memory BW. Although RT applied to MATH500 generates more tokens than Base, it still generates short enough sequences to under-utilize GPU memory BW. As a result, even at high batch sizes, the KV-cache remains manageable, and throughput scales proportionally with batch size increases.
However, RT applied to the LiveCodeBench exhibit a notable saturation in throughput at a batch size of approximately 128. On average, RT generates around 6588 output tokens for LiveCodeBench, shifting inference into a predominantly decode-bound stage. During the decode stage, loading the large KV-cache that scales linearly with batch size from GPU memory becomes the critical performance bottleneck, eventually saturating available GPU memory BW.
@@
Impact of LLM Serving System Optimizations
The most relevant optimization in our study is prefix caching, which stores the KV-cache of a query and reuses it whenever the same prefix appears in later queries. As shown in Table 2, prefix caching offers little energy savings for RT, since RT does not use additional tokens during the prefill stage. MV shows a more nuanced picture as prefix caching does reduce energy consumption, but less than expected. In theory, the normalized energy consumption for MV should approach 1 if queries are dominated by the prefill stage. However, we find that MV generates at least one sample with an output sequence an order of magnitude longer than the other samples. This costly event shifts the bottleneck from prefill to decode, making the decode stage the primary driver of energy usage in the benchmarks we evaluate.
System optimizations targeting the decode stage such as speculative decoding (10.5555/3618408.3619203; 2024arXiv240115077L; 2023arXiv230201318C; 2023arXiv230207863K) benefits RT and MV with prefix caching more than Base. Speculative decoding reduces the cost of the decode stage for TTC and Base by the same factor (10.5555/3618408.3619203) by using a smaller model to generate multiple tokens and calling an original model only to verify the tokens. The overall benefit for MV and RT is larger than that for Base as the decode stage is more dominant than the prefill stage for the two TTC approaches. Another major proposal disaggregated serving (2024arXiv240109670Z), separates prefill and decode across different clusters. This approach allows systems to reclaim idle resources, but incurs overhead from transferring the KV-cache between clusters. Importantly, just like speculative decoding, disaggregated serving applies its costs and savings uniformly to TTC and Base, maintaining the trend that we report in this study.
Table 2: Energy consumption of MV and RT with and without the use of prefix caching normalized to Base. Normalized energy consumption for RT stays similar as no tokens were added during the prefill stage. Prefix caching saves energy consumption for MV but smaller than anticipated.
MV RT
Tasks Prefix Cache No Prefix Cache Prefix Cache No Prefix Cache
Math 1.91 2.28 2.11 2.37
Code 2.45 3.52 4.85 4.94
Common Sense 4.96 4.75 1.01 1.01
@@ 5 Conclusion and Discussion
Our results (Figure 4) demonstrate that smaller models with TTC can compete effectively against larger models, with distinct Pareto frontiers emerging for each task difficulty level. This insight supports the implementation of difficulty-aware evaluators (gpt5; 2024arXiv240407720S; park2024large; dutulescu2024hard; xu2024adaption) that dynamically route queries to the most energy-efficient model achieving optimal accuracy. Optimal energy consumption $E_{o}$ of an oracle evaluator and realistic energy consumption $E_{r}$ of an imperfect evaluator in practical settings employing a high-performing model as a fallback can be calculated as
where $p_{i}$ denotes the probability that the evaluator correctly predicts difficulty level $i$, $q$ represents the probability of incorrect predictions, and $e_{m_{D}}$ is the energy consumed by the largest, most accurate model. Compared against the baseline approach ($E_{max}=Ne_{m_{D}}$), our analysis shows significant potential energy savings. Systems with an oracle evaluator consume 1.75 kWh compared to baseline consumptions of 2.03 kWh and 2.20 kWh, for LiveCodeBench and MATH500, respectively.